OpenStudy (anonymous):

1. Brett rolls a number cube labeled from 1 to 6 and tosses a penny at the same time. What is the probability that he will roll an even number and toss heads?

OpenStudy (anonymous):

I think it is 3/12 i.e 1/4

OpenStudy (anonymous):

yes 1/4

OpenStudy (anonymous):

A. 0.25 B. 0.33 C. 0.5 D. 0.66

OpenStudy (anonymous):

A

OpenStudy (anonymous):

1/4 = 0.25 so A is correct option pls do nt forget to click GOOD ANSWER button :)

OpenStudy (anonymous):

Probability of getting an even number is 3/6 Probability of getting a head is 1/2 the probability that he will roll an even number and toss heads = (3/6)*(1/2) = 1/4 http://www.oojih.com for free study guide and free practice questions